# Differentiable Cyclic Causal Discovery Under Unmeasured Confounders

### v1

@@ B.1 Hutchinson trace estimator for computing log determinant of the Jacobian
The estimator above still has a major drawback: computing the $\mathrm{Tr}(\mathbf{J}_{\mathbf{U}_{k}\mathbf{g}})$ still requires $\mathcal{O}(d^{2})$ gradient calls to compute exactly. Fortunately, Hutchinson trace estimator [44] can be used to stochastically approximate the trace of the Jacobian matrix. This then results in the following estimator that can be computed efficiently via reverse-mode automatic differentiation
@@ B.2 Parameter update via score maximization
As described in Section 3, the model parameters are updated in two stages. In the first stage, the parameters of the neural networks and the Gumbel-Softmax distribution, used to sample adjacency matrices, are updated via backpropagation using stochastic gradient descent. Since (5) forms an implicit block of an implicit normalizing flow, following, we directly estimate the gradients of $\hat{\mathcal{S}}(\mathbf{B})$ with respect to $\bm{x}$ and $\bm{\phi}=(\bm{\theta},\mathbf{B})$. The gradient computation involves two terms: $\frac{\partial}{\partial(\cdot)}\log\det(\mathbf{I}+\mathbf{J}_{\mathbf{g}_{x}}(\bm{x},\bm{\phi}))$ and $\frac{\partial\hat{\mathcal{S}}}{\partial\bm{z}}\frac{\partial\bm{z}}{\partial(\cdot)}$, where $(\cdot)$ is a placeholder for $\bm{x}$ and $\bm{\phi}$. From [38, 40], we use the following unbiased estimators for the gradients:
@@
where $\mathbf{G}_{z}(\bm{z})=\mathbf{g}_{z}(\bm{z},\bm{\phi})+\bm{z}$, and recall that $\mathbf{G}(\bm{x},\bm{z},\bm{\phi})=\mathbf{g}_{x}(\bm{x},\bm{\phi})+\bm{x}+\mathbf{g}_{z}(\bm{z},\bm{\phi})+\bm{z}$. See [38] for more details. The procedure SGUPDATE shown in Algorithm 1 performs the gradient computation in (35) and (36).
@@ B.3 DCCD-CONF and the baselines code details
We implemented our framework using the libraries Pytorch and Scikit-learn in Python and the code used in running the experiments can be found in the codes folder within the supplementary materials. We plan to make the code publicly available on GitHub upon publication of the paper.
Starting with an initialization of the model parameters $(\bm{\theta}^{(0)},\mathbf{B}^{(0)},\mathbf{\Sigma}_{Z}^{(0)})$, we iteratively alternate between maximizing the score function with respect to $(\bm{\theta}^{(t)},\mathbf{B}^{(t)})$ and $\mathbf{\Sigma}_{Z}^{(t)}$, as described in Algorithm 1. Standard stochastic gradient updates are used for $(\bm{\theta}^{(t)},\mathbf{B}^{(t)})$, while coordinate gradient descent, implemented via the Scikit-learn library, is applied to $\mathbf{\Sigma}_{Z}^{(t)}$. For modeling the causal function $\mathbf{g}_{x}$, we follow the setup of Sethuraman et al. [27], employing neural networks (NNs) with dependency masks parameterized by a Gumbel-softmax distribution. The log-determinant of the Jacobian is computed using a power series expansion combined with the Hutchinson trace estimator. To mitigate bias from truncating the power series expansion, the number of terms is sampled from a Poisson distribution, as detailed in Section 3.3 and Appendix B.1. The final objective is optimized using the Adam optimizer [56].
The learning rate in all our experiments was set to $10^{-2}$. The neural network models used in our experiments contained one multi-layer perceptron layer. No nonlinearities were added to the neural networks for the linear SEM experiments. We used tanh activation for the nonlinear SEM experiments and for the experiments on the perturb-CITE-seq data set. The graph sparsity regularization constant $\lambda$ was set to $10^{-3}$ for all the experiments. The sparsity inducing regularization constant for the inverse covariance matrix of the confounder distribution, $\rho$, was set to $10^{-2}$ in all the experiments. The models were trained and evaluated on NVIDIA RTX6000 GPUs.
@@
For NODAGS-Flow, we used the code provided by authors [27] available at https://github.com/Genentech/nodags-flows. The default values were set for the hyperparameters. We implemented the LLC algorithm based on the details provided in [26]. The implementation can be found within the codes/baselines folder in the supplementary materials. For FCI, we used the implementation that is available in the causallearn python library (https://github.com/py-why/causal-learn). For DCDI, we used the codebase provided by the authors [15], available at https://github.com/slachapelle/dcdi. The default hyperparameters were used while training and evaluating the model.
@@ B.4.1 Synthetic Experiments
where $\mathbf{W}$ is the weighted adjacency matrix.
@@
In this experiment, the number of observed nodes in the graph is fixed at d = 10 . Training data consists of combination of observational data and single-node interventions over all nodes, i.e., $\mathcal{I}=\emptyset\cup\{\{i\}\mid i\in[d]\}$, with 1000 samples per interventional setting. The confounder ratio (number of confounders divided by the number of nodes) is varied from 0.2 to 0.8.
@@
Here, we fix the confounder ratio at 0.4. The total number of nodes in the graph is varied from 10 to 80. As in the previous setup, training data consists of combination of observational data and single-node interventions across all nodes, with 1000 samples per interventional setting.
@@
We fix $d=10$ and set the confounder ratio to 0.4. The number of interventions during training varies from 0 to $d$. Zero interventions corresponds to observational data. When fewer than $d$ interventions are provided, the intervened nodes are selected arbitrarily. Each interventional setting consists of 1000 samples.
@@
In this case, we explicitly enforce a non-contractive causal mechanism $\mathbf{F}$ by rescaling edge weights to ensure that the Lipschitz constant of the edge weight matrix exceeds one. We set $d=10$ and provide observational data and single-node interventions across all nodes, with 1000 samples per intervention. The confounder ratio varies between 0.2 and 0.8.
@@ B.4.2 Gene Perturbation Data set
Table 2: The list of chosen genes from Perturb-CITE-seq dataset [48].
@@ C.3 Comparison with FCI-JCI
We also compare the performance of DCCD-CONF with FCI-JCI [29], which is an extension of FCI algorithm that is capable of handling multiple contexts (in this case interventional settings). FCI-JCI outputs a Partial Ancestral Graph (PAG), which is a graph structure that represents the equivalence class of MAGs. We define a modified SHD score in order to check if the DMG estimated by DCCD-CONF belongs to the same equivalence class of the ground truth DMG. To that end, we convert the ground-truth DMG and the estimated DMG to their augmented DMGs and then construct the MAG of the acyclified version of the augmented DMGs. The modified SHD score then computes the discrepancies in the conditions of Theorem A.19, i.e., we count: (i) the number of extra edges ($N_{1}$) using the skeletons of the estimated MAG and ground truth MAG, (2) number of mismatched unshielded colliders ($N_{2}$), and (3) discrepancies in the discriminating paths ($N_{3}$). Similarly, for FCI-JCI, we count the disagreements between the ground-truth MAG and the estimated PAG, i.e., mismatch in skeleton $(N_{1})$, mismatch in unshielded colliders ($N_{2}$), invariant edge orientation discrepancies ($N_{3}$). Finally, the modified SHD $N_{1}+N_{2}+N_{3}$. We compare DCCD-CONF and JCI-FCI over two different settings: (i) the training data consists of observational data and single node interventions over all the nodes in the graph, and (ii) the training data consists of observational data and interventions over 2 nodes (randomly chosen) in the graph. Due to the complexity of computing the modified SHD (as it involves iterating over the discriminating paths and inducing paths) we fix the number of observed nodes to be $d=5$. In the both the cases, the confounder ratio is varied between 0.2 and 0.8, and nonlinear SEM is used to generate the data. The results are summarized in Figure 8.

### v2

@@ B.1 Hutchinson trace estimator for computing log determinant of the Jacobian
The estimator above still has a major drawback: computing the $\mathrm{Tr}(\mathbf{J}_{\mathbf{U}_{k}\mathbf{g}})$ still requires $\mathcal{O}(d^{2})$ gradient calls to compute exactly. Fortunately, Hutchinson trace estimator [Hutchinson, 1989] can be used to stochastically approximate the trace of the Jacobian matrix. This then results in the following estimator that can be computed efficiently via reverse-mode automatic differentiation
@@ B.2 Parameter update via score maximization
As described in Section 3, the model parameters are updated in two stages. In the first stage, the parameters of the neural networks and the Gumbel-Softmax distribution, used to sample adjacency matrices, are updated via backpropagation using stochastic gradient descent. Since (5) forms an implicit block of an implicit normalizing flow, following Lu et al. [2021], we directly estimate the gradients of $\hat{\mathcal{S}}(\mathbf{B})$ with respect to $\bm{x}$ and $\bm{\phi}=(\bm{\theta},\mathbf{B})$. The gradient computation involves two terms: $\frac{\partial}{\partial(\cdot)}\log\det(\mathbf{I}+\mathbf{J}_{\mathbf{g}_{x}}(\bm{x},\bm{\phi}))$ and $\frac{\partial\hat{\mathcal{S}}}{\partial\bm{z}}\frac{\partial\bm{z}}{\partial(\cdot)}$, where $(\cdot)$ is a placeholder for $\bm{x}$ and $\bm{\phi}$. From [Lu et al., 2021, Behrmann et al., 2019], we use the following unbiased estimators for the gradients:
@@
where $\mathbf{G}_{z}(\bm{z})=\mathbf{g}_{z}(\bm{z},\bm{\phi})+\bm{z}$, and recall that $\mathbf{G}(\bm{x},\bm{z},\bm{\phi})=\mathbf{g}_{x}(\bm{x},\bm{\phi})+\bm{x}+\mathbf{g}_{z}(\bm{z},\bm{\phi})+\bm{z}$. See Lu et al. [2021] for more details. The procedure SGUPDATE shown in Algorithm 1 performs the gradient computation in (35) and (36).
@@ B.3 DCCD-CONF and the baselines code details
We implemented our framework using the libraries Pytorch and Scikit-learn in Python and the code used in running the experiments can be found in the following Github repository: https://github.com/muralikgs/dccd_conf.
Starting with an initialization of the model parameters $(\bm{\theta}^{(0)},\mathbf{B}^{(0)},\mathbf{\Sigma}_{Z}^{(0)})$, we iteratively alternate between maximizing the score function with respect to $(\bm{\theta}^{(t)},\mathbf{B}^{(t)})$ and $\mathbf{\Sigma}_{Z}^{(t)}$, as described in Algorithm 1. Standard stochastic gradient updates are used for $(\bm{\theta}^{(t)},\mathbf{B}^{(t)})$, while coordinate gradient descent, implemented via the Scikit-learn library, is applied to $\mathbf{\Sigma}_{Z}^{(t)}$. For modeling the causal function $\mathbf{g}_{x}$, we follow the setup of Sethuraman et al. [2023], employing neural networks (NNs) with dependency masks parameterized by a Gumbel-softmax distribution. The log-determinant of the Jacobian is computed using a power series expansion combined with the Hutchinson trace estimator. To mitigate bias from truncating the power series expansion, the number of terms is sampled from a Poisson distribution, as detailed in Section 3.3 and Appendix B.1. The final objective is optimized using the Adam optimizer [Kingma and Ba, 2015].
The learning rate in all our experiments was set to $10^{-2}$. The neural network models used in our experiments contained one multi-layer perceptron layer. No nonlinearities were added to the neural networks for the linear SEM experiments. We used tanh activation for the nonlinear SEM experiments and for the experiments on the perturb-CITE-seq data set. The graph sparsity regularization constant $\lambda$ was set to $10^{-2}$ for all the experiments. The sparsity inducing regularization constant for the inverse covariance matrix of the confounder distribution, $\rho$, was set to $10^{-1}$ in all the experiments. The models were trained and evaluated on NVIDIA RTX6000 GPUs.
@@
For NODAGS-Flow, we used the code provided by authors [Sethuraman et al., 2023] available at https://github.com/Genentech/nodags-flows. The default values were set for the hyperparameters. We implemented the LLC algorithm based on the details provided in [Hyttinen et al., 2012]. The implementation can be found within the codes/baselines folder in the supplementary materials. For FCI, we used the implementation that is available in the causallearn python library (https://github.com/py-why/causal-learn). For DCDI, we used the codebase provided by the authors [Brouillard et al., 2020], available at https://github.com/slachapelle/dcdi. The default hyperparameters were used while training and evaluating the model. For DAGMA and ADMG, we used the codebase provided by the authors Bello et al. [2022] and  Bhattacharya et al. [2021], available at https://github.com/kevinsbello/dagma and https://gitlab.com/rbhatta8/dcd respectively. We implemented LiNGAM-MMI based on the details provided by the authors [Suzuki and Yang, 2024].
@@ B.4.1 Synthetic Experiments
where $\mathbf{W}$ is the weighted adjacency matrix. In all the experiments, the training data consisted of 500 samples per interventional setting (unless specified otherwise).
@@
In this experiment, the number of observed nodes in the graph is fixed at d = 10 . Training data consists of combination of observational data and single-node interventions over all nodes, i.e., $\mathcal{I}=\emptyset\cup\{\{i\}\mid i\in[d]\}$. The confounder ratio (number of confounders divided by the number of nodes) is varied from 0.2 to 0.8.
Impact of Cycles
We fix the number of nodes in the graph $d=10$, The number of cycles in the graph is varied between 0 and 8, with the confounder ratio set to 0.3. The training data consists of observational data as well as single node interventions over all the nodes in the graph.
Impact of Nonlinearity
In this setting, the degree of nonlinearity (controlled by $\beta$) is varied between 0 and 1. That is, the data is generated from the following SEM:
$\bm{x}=(1-\beta)(\mathbf{W}^{\top}\bm{x}+\bm{z})+\beta\tanh(\mathbf{W}^{\top}\bm{x}+\bm{z}).$
Here, $\beta=0$ implies the SEM is purely linear and $\beta=1.0$ implies the data is purely nonlinear. The confounder ratio is set to 0.3. The number of cycles is randomly set.
@@
We fix the confounder ratio at 0.4. The total number of nodes in the graph is varied from 10 to 80. As in the previous setup, training data consists of combination of observational data and single-node interventions across all nodes, with 500 samples per interventional setting.
@@
We fix $d=10$ and set the confounder ratio to 0.4. The number of interventions during training varies from 0 to $d$. Zero interventions corresponds to observational data. When fewer than $d$ interventions are provided, the intervened nodes are selected arbitrarily. Each interventional setting consists of 500 samples.
@@
In this case, we explicitly enforce a non-contractive causal mechanism $\mathbf{F}$ by rescaling edge weights to ensure that the Lipschitz constant of the edge weight matrix exceeds one. We set $d=10$ and provide observational data and single-node interventions across all nodes, with 500 samples per intervention. The confounder ratio varies between 0.2 and 0.8.
@@ B.4.2 Gene Perturbation Data set
Table 2: The list of chosen genes from Perturb-CITE-seq dataset [Frangieh et al., 2021].
@@ C.3 Comparison with FCI-JCI
Here, we compare the performance of DCCD-CONF with FCI-JCI Mooij et al. [2020], which is an extension of FCI algorithm that is capable of handling multiple contexts (in this case interventional settings). FCI-JCI outputs a Partial Ancestral Graph (PAG), which is a graph structure that represents the equivalence class of MAGs. We define a modified SHD score in order to check if the DMG estimated by DCCD-CONF belongs to the same equivalence class of the ground truth DMG. To that end, we convert the ground-truth DMG and the estimated DMG to their augmented DMGs and then construct the MAG of the acyclified version of the augmented DMGs. The modified SHD score then computes the discrepancies in the conditions of Theorem A.19, i.e., we count: (i) the number of extra edges ($N_{1}$) using the skeletons of the estimated MAG and ground truth MAG, (2) number of mismatched unshielded colliders ($N_{2}$), and (3) discrepancies in the discriminating paths ($N_{3}$). Similarly, for FCI-JCI, we count the disagreements between the ground-truth MAG and the estimated PAG, i.e., mismatch in skeleton $(N_{1})$, mismatch in unshielded colliders ($N_{2}$), invariant edge orientation discrepancies ($N_{3}$). Finally, the modified SHD $N_{1}+N_{2}+N_{3}$. We compare DCCD-CONF and JCI-FCI over two different settings: (i) the training data consists of observational data and single node interventions over all the nodes in the graph, and (ii) the training data consists of observational data and interventions over 2 nodes (randomly chosen) in the graph. Due to the complexity of computing the modified SHD (as it involves iterating over the discriminating paths and inducing paths) we fix the number of observed nodes to be $d=5$. In the both the cases, the confounder ratio is varied between 0.2 and 0.8, and nonlinear SEM is used to generate the data. The results are summarized in Figure 8.
Figure 9: Performance comparison between DCCD-CONF and LiNGAM-MMI on $d=5$ node graphs.
@@
C.4 Comparison with LiNGAM-MMI
We compare the performance of DCCD-CONF with LiNGAM-MMI Suzuki and Yang [2024], which extends the ICA-based LiNGAM Shimizu et al. [2006] to handle hidden confounding. Since LiNGAM-MMI requires iterating over all possible node permutations, its computational cost grows rapidly with the number of nodes; hence, we restrict our comparison to $d=5$. The training data include both observational samples and single-node interventions for all nodes, with 500 samples per interventional setting. We vary the confounder ratio (i.e., the ratio of confounders to observed nodes) between 0.2 and 0.8. As shown in Figure 9, DCCD-CONF consistently outperforms LiNGAM-MMI across all tested confounder ratios.
C.5 Hyperparameter Sensitivity
We evaluate the sensitivity of DCCD-CONF to its hyperparameters: (i) the directed edge sparsity regularization coefficient $\lambda_{c}$, and (ii) the bidirected edge sparsity regularization coefficient $\rho$. Figure 10 summarizes the results for $\lambda_{c},\rho\in{0.1,0.01,0.001}$ on graphs with $d=10$ nodes and a confounder ratio of 0.3. The model was trained using both observational data and all single-node interventions. As shown in the figure, DCCD-CONF remains fairly robust to hyperparameter variations, achieving normalized SHD values below 0.3 and F1 scores above 0.65 for most combinations of $\lambda_{c}$ and $\rho$.
Figure 10: Illustration of DCCD-CONF performance for various choices of hyperparameters $\lambda_{c}$ (directed edge sparsity regularization coeff.) and $\rho$ (bidirected edge sparsity regularization coeff.).
C.6 Training Time Comparison
Figure 11 compares the training times of DCCD-CONF and baseline methods. Unlike the other algorithms, LLC does not require stochastic gradient-based training, as it relies solely on solving a series of linear regressions. Consequently, LLC is considerably faster, as shown in Figure 11. Excluding LLC, NODAGS-Flow is the most efficient in terms of runtime; however, it cannot account for confounders within its framework. DCCD-CONF achieves training times comparable to ADMG and DCDI while simultaneously handling confounders, cycles, and nonlinear dependencies. The training times are computed on $d=10$ node graphs with training dataset consisting of observational and all single-node interventions. DCCD-CONF was trained for 200 epochs. All models were run on RTX6000 GPUs.
Figure 11: Training time comparison between DCCD-CONF and the baselines
C.7 Additional Results on Perturb-CITE-seq Dataset
In addition to test-set NLL, we evaluate the performance of DCCD-CONF and the baselines on the Perturb-CITE-seq dataset Frangieh et al. [2021] using Interventional Mean Absolute Error (I-MAE) as the evaluation metric. I-MAE is computed as the mean of $\|\bm{x}+\mathbf{g}_{x}(\bm{x})\|_{1}/d$ over all observations $\bm{x}$ in the held-out test set. Beyond the baselines discussed in Section 4, we also include Bicycle Rohbeck et al. [2024] and DCDFG Lopez et al. [2022] for comparison. Bicycle supports nonlinear and cyclic structures, whereas DCDFG restricts the search space to acyclic graphs. The results, summarized in Table 3, show that DCCD-CONF remains competitive with state-of-the-art methods under the I-MAE metric.
Table 3: Results on Perturb-CITE-seq Frangieh et al. [2021] gene perturbation dataset. The table presents the average Mean Absolute Error (MAE) on the test set, averaged over multiple trials (standard deviation is reported within paranthesis).
Method Control Co-Culture IFN-$\gamma$
DCCD-CONF 0.781 (0.037) 0.765 (0.046) 0.843 (0.035)
NODAGS 0.847 (0.018) 0.762 (0.018) 0.861 (0.023)
Bicycle 0.782 (0.042) 0.735 (0.036) 0.883 (0.028)
DCDFG 0.845 (0.066) 0.774 (0.038) 0.891 (0.041)
Additionally, we report the adjacency matrix of the recovered causal graph for the cell condition “Co-Culture” in Figure 12. DCCD-CONF identified 38 feedback cycles. This number validates prior work showing that gene regulatory networks are rich in feedback loops Frangieh et al. [2021].
Figure 12: Adjacency matrix of learnt by DCCD-CONF for “Co-Culture” cell condition of Perturb-CITE-seq dataset.
C.8 Protein Signaling Dataset
We further evaluate DCCD-CONF on a biological dataset for protein signaling network discovery Sachs et al. [2005], which is widely used as a benchmark for causal discovery algorithms.
Table 4: Performance comparison on Sachs et al. [2005] protein signaling dataset.
Method SHD
DCCD-CONF 18
DAG-GNN Yu et al. [2019] 19
DAGMA 21
NOTEARS 22
The dataset contains continuous measurements of multiple phosphorylated proteins and phospholipid components in human immune system cells, with the corresponding network capturing the ordering of interactions among pathway components. Based on $n=7466$ samples across $m=11$ cell types, Sachs et al. [2005] identified 20 edges in the underlying graph. Using the consensus network from Sachs et al. [2005] as ground truth, we evaluate performance using the Structural Hamming Distance (SHD) as the error metric. The results, summarized in Table 4, show that DCCD-CONF performs comparably to the baselines. The recovered directed graph is visualized in Figure 13.
Figure 13: (Left) Consensus graph from Sachs et al. [2005], (right) graph learned by DCCD-CONF.